Find the sum of four consecutive negative integers, such that the square of the sum of the first and fourth is 289. Thanks for your help.
Answer by ewatrrr(24785)

four consecutive negative integers: x, x+1, x+2, x+3
(2x+3)^2 = 289
4x^2 + 12x + 9 = 289
4x^2 + 12x -280 = 0
x^2 + 3x -70 = 0
(x+10)(x-7) = 0
x = -10
The four consecutive negative integers are: -10, -9, -8, -7
Their sum = -34